please help me solve this problem +please+help+me+solve+this+problem+ A number is increased by 2 thirds of the number is 35. find the number
Answer by bucky(2189)

Let N be the unknown number.
.
"Increased by 2 thirds of the number" means find 2 thirds of N and add it to (or increase) N by it.
.
So we have N and increase it by 2 thirds N, and this translates to:
.
N + (2/3)N
.
According to the problem, this equals 35. So we have an equation:
.
N + (2/3)N = 35
.
We can get rid of the denominator 3 by multiplying both sides (all terms) by 3 and we then have:
.
3N + 2N = 105
.
Adding together both terms on the left side reduces the equation to:
.
5N = 105
.
Solve for N by dividing both sides of this equation by 5 to get:
.
N = 105/5 = 21
.
That's the answer. The number is 21. 1 third of 21 is 7 so 2 thirds of 21 is 14. When you increase the number (21) by 2 thirds of 21 it is:
.
21 + 14
and this equals 35, just as the problem says it should.
.
Hope this helps you to understand this problem and how you can solve it.
.